OpenStudy (caitlyn):

Can someone help explain this process to me? Its on the Radical Expressions and Equations Unit Test. square root sign/250h^4k^5. I'm not really sure how you would write that, but I really need to understand the process of how to solve this.

OpenStudy (jdoe0001):

how to write what?

OpenStudy (caitlyn):

square root sign/250h^4k^5.

OpenStudy (jdoe0001):

do you just need to simplify it?

OpenStudy (caitlyn):

I only really know how to do really simple problems.

OpenStudy (caitlyn):

I think i need to actually solve it but simplifying is a start.

OpenStudy (jdoe0001):

\( 250 \implies 25*5*2 \)

OpenStudy (caitlyn):

okay. I understand that.

OpenStudy (jdoe0001):

and thus \( \sqrt{25}*\sqrt{5*2} \implies 5\sqrt{10}\)

OpenStudy (caitlyn):

But what do you do with the exponents?

OpenStudy (jdoe0001):

\( 250 \implies 25*5*2 \) and thus \( \sqrt{25}*\sqrt{5*2} \implies 5\sqrt{10}\) $$ h^4=\pmatrix{h^2}^2\\ k^5=\pmatrix{k^2}^2*k^1\\ \text{so what you have in the end is}\\ \sqrt{5^2*10*\pmatrix{h^2}^2*\pmatrix{k^2}^2*k} $$

OpenStudy (caitlyn):

Okay. I think i got that. Is that it? Like, would that be the final stage in solving the problem or is there another step?